OpenStudy (anonymous):

hey guys I need to solve this equation but my calculator hates me: x=14(tan 42degrees)

OpenStudy (nubeer):

12.6

OpenStudy (anonymous):

That's what I got, but these are the answer choices: 3.7 11.5 3.3 4.5

OpenStudy (anonymous):

tan(14) = 0.90040404429783994512047720388537 Multiply that by 14 and you get nubeer's answer.

OpenStudy (anonymous):

I did that, but its not one of the answer choices

OpenStudy (anonymous):

Okay, so there's something we don't know. Can you show us the source of your question?

OpenStudy (anonymous):

ya hold on

OpenStudy (nubeer):

if u r using tan then u have to use base value which is 5

OpenStudy (anonymous):

ok how to do that tho

OpenStudy (nubeer):

tan (theta) = x / 5 tan (42) *5 = x

OpenStudy (anonymous):

Okay, tan(angle) is opposite/adjacent So opposite = tan(angle)*adjacent. What "we didn't know" is that the adjacent length is "5". So your answer is x = tan(42)*5

OpenStudy (anonymous):

PS: I have no idea where that "14" came from...